# When do Reforms meet fairness concerns in school admissions?

### v3

@@ 3.1. Fairness by stability
In this example, we illustrate how the constrained $GS$ mechanism has shortened the chains needed to reach a stable matching. It is well known that this type of chains lead to unambiguous welfare losses: each student in the chain is worse off, all other students are unaffected (Kesten,, 2010). [^10] These chains are initiated by the so-called interrupters. These are students who initiate chains of rejections that return to them (Kesten,, 2010). However, under the Boston mechanism, (where all schools are first-preference-first schools) there is no such chain, and thus the constrained Boston mechanism can be compared to the constrained Gale-Shapley mechanism. We also show the comparison result for the Gale-Shapley mechanisms with different constraints.
@@
Suppose that there are at least two schools and let $k>1$.
@@
Consider a problem with $n$ students and $m$ schools such that for each $s,s^{\prime}\in S$, $\succ_{s}\mathrel{=}\succ_{s^{\prime}}$. We assume that students have tier preferences. The set $S$ of schools is partitioned into two sets $S_{1},S_{2}$. Each student $i$ prefers each school in $S_{1}$ to each school in $S_{2}$. [^11] Coles et al., (2013) observed that the academic job market has this structure and referred to it as block correlated preferences. We assume that each student finds each school acceptable and $n>\sum_{s}q_{s}$.
@@
While this example is simplified to illustrate the fact that fairness by stability might not be a relevant notion for comparing the mechanisms that we study, all that is necessary for the result is that many students rank a subset of schools such that they exhaust the listing constraint, and such that these schools do not have enough seats to accommodate all these students.
@@ 3.2. Fairness by counting
In this section we introduce a criterion of comparing the number of blocking students. With this criterion, mechanisms can be compared at problems where both induce unstable outcomes.
@@
The intuition of this result is that by shortening the schools listed by student $i_{2}$, she is worse off while the other students are weakly better off. First, she is a blocking student for the new matching. Second, student $i_{1}$ is not a blocking student for the new matching, though she was a blocking student for the old matching. But a new blocking student appears so that there are two blocking students in total.
This example turns out to be general pattern. When students shorten the list, the set of blocking students changes, but the size of this set never decreases. By sequentially applying this argument to all students, we get the following result. [^12] We are very grateful to a referee for suggesting a very simplified technique for proving this result.
@@
Consider the Chinese mechanisms $Ch^{(1)}=\beta$ and $Ch^{(3)}$ and note that for the problem $(P,\succ,q)$ specified in that example, $Ch^{(1)}(P,\succ,q)=\beta^{3}(P,\succ,q)$ and $Ch^{(3)}(P,\succ,q)=GS^{3}(P,\succ,q)$. According to the conclusion in Example 4, there are more blocking students for $Ch^{(3)}(P,\succ,q)$ than $Ch^{(1)}(P,\succ,q)$. According to Chen and Kesten, (2017), there is a problem where $Ch^{(3)}$ produces a stable outcome but $Ch^{(1)}$ does not.
@@
Doğan and Ehlers, (2021) compare mechanisms by the inclusion of blocking pairs and blocking students. However, these criteria are stronger than fairness by counting (if the set of blocking pairs or blocking students shrinks, then the number of blocking students does as well) and will lead to negative results for our comparisons. To see this, consider Example 4. In this example, $(i_{5},s_{5})$ is a blocking pair for $SD^{4}(P,\succ,q)$ but not for $\beta^{4}(P,\succ,q)$. In addition, $(i_{4},s_{4})$ is a blocking pair for $\beta^{4}(P,\succ,q)$ but not for $SD^{4}(P,\succ,q)$.
@@
4. Fairness in strategic settings
We study mechanisms when students report their preferences truthfully. This approach is standard, but it has a limited interpretation since students may have incentives to misreport their preferences in the mechanisms in question. We face the following methodological difficulty. On the one hand, when students sincerely report their true preferences, the set of blocking students is well understood but does not necessarily represent the blocking students of the expected outcomes. On the other hand, when students are well sophisticated, potential blocking students may misrepresent their preferences such that the expected outcome does not exhibit any blocking pair. For example, in a game induced by the Boston mechanism, all Nash equilibrium outcomes are stable (Ergin and Sönmez,, 2006; Haeringer and Klijn,, 2009).
To address this difficulty, we introduce a new method of studying fairness. The general principal is that some students are not well sophisticated to best respond and thus report truthfully; and we study whether the mechanisms are fair to these students. We assume that some students are sincere while others are sophisticated. We check if some results from the previous section are robust to this analysis. Note that all negative results remain true in this setting.
Reforms in Chicago. We study some reforms in Chicago. The Chicago selective high school system called for a reform in the middle of its admissions process in 2009 and initiated another reform in 2010. The admission to each of these schools is very competitive. Each school uses a common priority based on students’ composite scores. Prior to 2009, the school board was using the Boston mechanism with a ranking constraint of 4 schools. In 2009, they replaced this mechanism with the serial dictatorship with the same ranking constraint. A year after, they maintained the serial dictatorship rule but extended the ranking constraint to 6 schools. Their motivation was that some students couldn’t figure out the optimal play so as to avoid losing their high priorities.
Following Pathak and Sönmez, (2008), we consider a model with sincere students and sophisticated students. Let $N$ denote the set of sincere students and $M$ the set of sophisticated students such that $N\cup M=I$ and $N\cap M=\emptyset$. Sincere students always report their preferences truthfully while sophisticated students may strategically misreport their preferences. The idea is to measure unfairness with respect to the set of blocking students among sincere students.
Any mechanism $\varphi$ and a pair $(\succ,q)$ induce a normal form game such that students are the players, the strategies are the preference reports, and the outcome function is $\varphi(.,\succ,q)$. Let $(\varphi,P)$ denote the game induced by $\varphi$. A strategy profile $P^{\prime}=(P_{N},P^{\prime}_{M})$ is a Nash equilibrium of the game $(\varphi,P)$ if for each sophisticated student $i\in M$, there is no strategy $P^{\prime\prime}_{i}$ such that $\varphi_{i}(P^{\prime\prime}_{i},P^{\prime}_{-i},\succ,q)\mathrel{P_{i}}\varphi_{i}(P^{\prime},\succ,q)$.
We consider the 2010 reform. Recall that the school board has maintained the serial dictatorship mechanism but extended the ranking constraint from 4 to 6 schools. The game under the constrained serial dictatorship mechanism also has a unique Nash equilibrium outcome.
Lemma 1.
Let $k>1$ and suppose that there are sincere students and sophisticated students. For any problem where schools have a common priority, the constrained serial dictatorship mechanism $SD^{k}$ has a unique Nash equilibrium outcome (where sincere students play truthfully and sophisticated students best respond). The equilibrium outcome is the matching $SD(P^{k}_{N},P_{M},\succ,q)$.
The proof of this result is constructive. It consists of showing that the matching $SD(P^{k}_{N},P_{M},\succ,q)$, where the constraint only applies to sincere students, is the unique Nash equilibrium outcome of the game. The reason is that this is the unique stable matching under $(P^{k}_{N},P_{M},\succ,q)$ and every sophisticated student who is part of a blocking pair with a school can obtain a seat at this school by ranking it first and picking it at her turn.
Proof.
Let $(P_{N},P^{\prime}_{M})$ be a Nash equilibrium of the game $(SD^{k},P)$ where sincere students report truthfully $P_{N}$ and sophisticated students report $P^{\prime}_{M}$.
We claim that $SD^{k}(P_{N},P^{\prime}_{M},\succ,q)$ is stable at $(P^{k}_{N},P_{M},\succ,q)$. Clearly, under the constrained profile, there is no blocking pair involving a sincere student. Suppose that there is a blocking pair $(i,s)$ where $i$ is a sophisticated student. Then at least one seat of $s$ is either unassigned or has been assigned to a student with lower ranking according to the order in which students pick schools. Then at $i$’s turn, school $s$ still has a seat available. Let $P_{i}^{s}$ be a preference relation where $i$ has ranked school $s$ first. Then at her turn, student $i$ will pick school $s$ under $(P^{k}_{N},P^{s}_{i},P^{\prime}_{-(N\cup\{i\})})$. This contradicts the assumption that $(P_{N},P^{\prime}_{M})$ is a Nash equilibrium of the game $(SD^{k},P)$.
Next $SD^{k}(P_{N},P^{\prime}_{M},\succ,q)$ is individually rational under $P$. Clearly, it is individually rational to each sincere student. If it is not individually rational to a sophisticated student $i$ under $P_{i}$, then she has ranked under $P^{\prime}_{i}$ a school $s$ that is unacceptable under $P_{i}$ and to which she is matched to under $SD^{k}(P_{N},P^{\prime}_{M},\succ,q)$. Student $i$ is better reporting a preference relation where she finds no school acceptable. This contradicts the fact that $(P_{N},P^{\prime}_{M})$ is a Nash equilibrium of the game $(SD^{k},P)$.
Since schools have the same priority order, there is a unique stable matching under $(P^{k}_{N},P_{M},\succ,q)$, and thus the Nash equilibrium is unique. The equilibrium outcome is the matching $SD(P^{k}_{N},P_{M},\succ,q)$. ∎
The game under the constrained serial dictatorship mechanism can be interpreted as if the constraint applies only to sincere students. All together, we obtain the following results.
Theorem 2.
Let $k>\ell>1$. Suppose that there are sincere students who play truthfully and sophisticated students who best respond. For any problem where schools have a common priority, the Nash equilibrium outcome of the constrained serial dictatorship mechanism $SD^{\ell}$ has at least as many blocking students as the Nash equilibrium outcome of the constrained serial dictatorship mechanism $SD^{k}$.
@@ 5. Stability and manipulability
In this section we will elucidate the relation between blocking students and manipulating students, i.e., those who may benefit from misrepresenting their preferences to the mechanisms.
@@
It turns out that there is a relation between blocking students and manipulating students for the constrained Boston mechanism and the constrained Gale-Shapley mechanism. Interestingly, these relations for the two mechanisms are reversed.
@@
Note that there are instances where the reverse of parts (i) and (ii) of the theorem are not true. For part (ii) see Example 5. For part (i) consider Example 4 and suppose that there, student $i_{n}$ prefers school $s_{1}$ first, $s_{4}$ next and $s_{5}$ last. The preferences of the remaining students are unchanged. Then under $\beta^{3}$ student $i_{n}$ is matched to school $s_{5}$ and is not a blocking student. However by ranking school $s_{4}$ first as in the example she is matched to it. That is she is manipulating student of $\beta^{3}$.
@@
The manipulation strategy under the constrained GS is to include an unlisted acceptable school in the list. But when the constrained GS is stable, all the seats of such a school are assigned to higher priority students, and such a manipulation does not help. This implies that constrained the serial dictatorship mechanism is non-manipulable and stable for the same set of problems.
@@
This matching is not stable at $(P,\succ,q)$ because student $i_{2}$ is unmatched, finds school $s_{3}$ acceptable while student $i_{4}$ is matched to it and $i_{2}\mathrel{\succ_{s_{3}}}i_{4}$. We claim that $GS^{2}$ is not manipulable at $(P,\succ,q)$. Only student $i_{2}$ could benefit from misrepresenting her preferences to the mechanism $GS^{2}$ because each of the other students is matched to her most-preferred school. Let $P_{i_{2}}^{s_{3}}$ be a preference relation where student $i_{2}$ has ranked only school $s_{3}$ acceptable. Then,
@@ 6. Conclusions
In response to various concerns, many school districts around the world have recently reformed their admissions systems. The reforms concern essentially two major changes. First, they replaced the immediate acceptance procedure ( Boston mechanism) with the Gale-Shapley’s student-proposing deferred acceptance procedure while maintaining ranking constraints. Second, some school districts kept using the Gale-Shapley mechanism but extended the number of schools that each student is allowed to report. Anecdotal evidence points to both the vulnerability to manipulation and fairness as reasons for these reforms. We showed that the immediate acceptance procedure has weakly more manipulating students than blocking students and the reverse for the Gale-Shapley’s. We demonstrated that extending ranking constraints in the Gale-Shapley mechanism led to fewer number of blocking students.
The fact that constrained Gale Shapley has relatively more blocking students (than it has manipulating students) suggests that in theory fairness of this mechanism might be of a stronger concern than manipulability. Simultaneously, relaxing the constraint in this mechanism is guaranteed to decrease the number of blocking students. Whether this theoretical coincidence had any practical relevance for the reforms under consideration is an open question and requires further empirical research.
@@
To simplify the exposition we divide the appendix in three subsections. In each subsection, we order the results in logical order. All mechanisms that we consider are individually rational. We only consider blocking pairs to check for (the violation of) stability. We first present a useful lemma.
@@
Appendix A: Proof of Proposition 1, 3
@@
Part (ii): Suppose that $\mu=FPF^{k}(P,\succ,q)$ is stable at $(P,\succ,q)$. By Claim 2 every matched student under $GS^{k}(P,\succ,q)$ is not a manipulating student of $GS^{k}$ at $(P,\succ,q)$. It is enough to show that no unmatched student under $GS^{k}(P,\succ,q)$ has a profitable misrepresentation. Because $GS^{k}$ is individually rational, by Claim 1, we further need to restrict ourselves to manipulation by top ranking schools. Since $\mu$ is stable at $(P,\succ,q)$, we claim that it is also stable at $(P^{k},\succ,q)$. Since $GS^{k}$ is individually rational, we need to check that there is no blocking pair. Suppose, to the contrary, that a pair $(i,s)$ is a blocking pair for $\mu$ under $(P^{k},\succ,q)$. Then, $s\mathrel{P_{i}^{k}}\mu(i)$ and either (i) school $s$ has an empty seat under $\mu$ or (ii) there is a student $j$ such that $\mu(j)=s$ and $i\mathrel{\succ_{s}}j$. Note that $s\mathrel{P^{k}_{i}}\mu(i)$ implies that $s\mathrel{P_{i}}\mu(i)$. Therefore, $(i,s)$ is also a blocking pair for $\mu$ under $(P,\succ,q)$, thus contradicting our assumption that $\mu$ is stable at $(P,\succ,q)$. Therefore $\mu$ is stable at $(P^{k},\succ,q)$. Since $GS(P^{k},\succ,q)$ is the student-optimal stable matching under $(P^{k},\succ,q)$,
@@
Part (i): The Boston mechanism is a special case of the First-Preference-First mechanism when every school is a first-preference-first school. Suppose that $\beta^{k}(P,\succ,q)$ is stable at $(P,\succ,q)$. By equation 1, each student finds the outcome $GS^{k}(P,\succ,q)$ at least as good as $\beta^{k}(P,\succ,q)$ under $P^{k}$. We also know that the Boston mechanism is Pareto efficient, that is, for each problem there is no other matching that each student finds at least as good as its outcome (Abdulkadiroğlu and Sönmez,, 2003). Therefore, the matching $\beta^{k}(P,\succ,q)=\beta(P^{k},\succ,q)$ is Pareto efficient under $P^{k}$. Thus, $GS^{k}(P,\succ,q)=\beta^{k}(P,\succ,q)$ is stable at $(P,\succ,q)$.
@@
Appendix B: Proof of Theorems 1, 2
@@
We prove the it by the contradiction. Suppose, to the contrary, that student $i$ is a blocking student for $\mu$ under $(P,\succ,q)$ such that $\mu(i)=s$ for some school $s$. Then, there is a school $s^{\prime}$ such that $s^{\prime}\mathrel{P_{i}}\mu(i)$ and either (i) $\lvert\mu^{-1}(s^{\prime})\lvert<q_{s^{\prime}}$ or (ii) there is a student $j$ such that $\mu(j)=s^{\prime}$ and $i\mathrel{\succ_{s^{\prime}}}j$. Since $\mu(i)=s$, school $s$ is one of the top $x$ acceptable schools under $P_{i}$ where $x=\ell$ if $i\in N$ and $x=k$ if $x\notin N$. Thus $s^{\prime}\mathrel{P^{x}_{i}}\mu(i)=s$ and $(i,s^{\prime})$ is a blocking pair of $\mu$ under $(P^{\ell}_{N},P^{k}_{-N},\succ,q)$, contradicting the stability of $\mu$ under $(P^{\ell}_{N},P^{k}_{-N},\succ,q)$.
@@
We call on to the sequential version of McVitie and Wilson, (1970) of the deferred acceptance algorithm. This is a version where students apply one at a time according to a predetermined order such that in each step the highest ordered student among the ones whose applications have not yet been tentatively accepted applies.
The idea of the proof is to consider close ranking constraints $k-1$ and $k$, where $k>1$, and replacing students’ preference relations in $P^{k-1}$ by the ones in $P^{k}$. Let $N\subsetneq I$ be a proper subset of $I$ and $i\notin N$. Suppose that starting from $P^{k-1}$ we have replaced all the preferences of students in $N$ by their preferences in $P^{k}$ and define $\widehat{P}=(P_{N}^{k},P^{k-1}_{-N})$. Note that $N$ may be empty. Next we consider student $i$: from $\widehat{P}$ we replace her preference relation $P^{k-1}_{i}$ by $P^{k}_{i}$. Let $\nu=GS(P_{i}^{k-1},\widehat{P}_{-i},\succ,q)$ and $\mu=GS(P^{k}_{i},\widehat{P}_{-i},\succ,q)$.

### v4

@@ 3.1. Fairness by stability
This example illustrates how the constrained $GS$ mechanism has shortened the chains needed to reach a stable matching. It is well known that this type of chain leads to unambiguous welfare losses. Each student in the chain is worse off, and all other students are unaffected (Kesten,, 2010). [^9] These chains are initiated by the so-called interrupters. These are students who initiate chains of rejections that return to them (Kesten,, 2010). However, under the Boston mechanism, (where all schools are first-preference-first schools) there is no such chain, and thus the constrained Boston mechanism can be compared to the constrained Gale-Shapley mechanism. We also show the comparison result for the Gale-Shapley mechanisms with different constraints.
@@
Let $\ell$ and $k$ be positive integers such that $k>\ell>1$ and suppose that there are at least two schools:
@@
Chen and Kesten, (2017) have established that any (unconstrained) Chinese mechanism $Ch^{e}$ is more stable than any Chinese mechanism $Ch^{e^{\prime}}$ where $e^{\prime}=ke$ for $k\in\mathbb{N}\cup\{\infty\}$. Their result and ours are similar but not corollary of each other. Indeed, $Ch^{1}$ is the Boston mechanism and $Ch^{\infty}$ is the Gale-Shapley mechanism such that for a problem $(P,\succ,q)$, we can write $\beta^{k}(P,\succ,q)=Ch^{(1)}(P^{k},\succ,q)$ and $GS^{k}(P,\succ,q)=Ch^{\infty}(P^{k},\succ,q)$. Our results concern constrained Chinese mechanisms where both the parameter $e$ and the constraint $k$ could be a source of blocking while in Chen and Kesten, (2017), the parameter $e$ is the only source of blocking.
@@
Consider a problem with $n$ students and $m$ schools such that for each $s,s^{\prime}\in S$, $\succ_{s}\mathrel{=}\succ_{s^{\prime}}$. We assume that students have tier preferences. The set $S$ of schools is partitioned into two sets $S_{1},S_{2}$. Each student $i$ prefers each school in $S_{1}$ to each school in $S_{2}$. [^10] Coles et al., (2013) observed that the academic job market has this structure and referred to it as block-correlated preferences. We assume that each student finds each school acceptable and $n>\sum_{s}q_{s}$.
@@
But we can still distinguish different constrained Gale-Shapley mechanisms by the number of blocking students within this preference domain. For example, suppose that students have the same ranking over all schools in $S_{1}$ and we compare $GS^{k-1}$ and $GS^{k}$. Every unmatched student is a blocking student (and vice versa), and $GS^{k}$ matches strictly more students than $GS^{k-1}$ and thus $GS^{k}$ has fewer blocking students than $GS^{k-1}$ does.
While this example is simplified to illustrate the fact that fairness by stability might not be able to distinguish the mechanisms that we study in real-life instances, all that is necessary for the result is that many students rank a subset of schools such that they exhaust the listing constraint, and such that these schools do not have enough seats to accommodate all these students.
@@ 3.2. Fairness by counting
In this section, we introduce a criterion for comparing the number of blocking students. With this criterion, mechanisms can be compared at problems where both induce unstable outcomes.
@@
The intuition of this result is that by shortening the schools listed by student $i_{2}$, she is worse off while the other students are weakly better off. First, she is a blocking student for the new matching. Second, student $i_{1}$ is not a blocking student for the new matching, though she was a blocking student for the old matching. But a new blocking student appears so there are two blocking students in total.
This example turns out to be a general pattern. When students shorten the list, the set of blocking students changes, but the size of this set never decreases. By sequentially applying this argument to all students, we get the following result. [^11] We are very grateful to a referee for suggesting a much simplified technique for proving this result.
@@
Consider the Chinese mechanisms $Ch^{(1)}=\beta$ and $Ch^{(3)}$ and note that for the problem $(P,\succ,q)$ specified in Example 4, $Ch^{(1)}(P,\succ,q)=\beta^{3}(P,\succ,q)$ and $Ch^{(3)}(P,\succ,q)=GS^{3}(P,\succ,q)$. According to the conclusion in Example 4, there are more blocking students for $Ch^{(3)}(P,\succ,q)$ than $Ch^{(1)}(P,\succ,q)$. According to Chen and Kesten, (2017), there is a problem where $Ch^{(3)}$ produces a stable outcome but $Ch^{(1)}$ does not.
@@
Two other notions, comparing mechanisms by the inclusion of blocking pairs and blocking students, have also been studied by Doğan and Ehlers, (2021). However, these criteria are stronger than fairness by counting (if the set of blocking pairs or blocking students shrinks, then the number of blocking students does as well) and will lead to negative results for our comparisons. To see this, consider Example 4. In this example, $(i_{5},s_{5})$ is a blocking pair for $SD^{4}(P,\succ,q)$ but not for $\beta^{4}(P,\succ,q)$. In addition, $(i_{4},s_{4})$ is a blocking pair for $\beta^{4}(P,\succ,q)$ but not for $SD^{4}(P,\succ,q)$.
@@ 4. Stability and manipulability
In this section, we will elucidate the relation between blocking students and manipulating students, i.e., those who may benefit from misrepresenting their preferences to the mechanisms.
@@
It turns out that there is a relationship between blocking students and manipulating students for the constrained Boston mechanism and the constrained Gale-Shapley mechanism. Interestingly, these relations between the two mechanisms are reversed.
@@
Note that there are instances where the reverse of parts (i) and (ii) of the theorem are not true. For part (ii) see Example 5. For part (i) consider Example 4 and suppose that there, student $i_{n}$ prefers school $s_{1}$ first, $s_{4}$ next, and $s_{5}$ last. The preferences of the remaining students are unchanged. Then under $\beta^{3}$ student $i_{n}$ is matched to school $s_{5}$ and is not a blocking student. However, by ranking school $s_{4}$ first as in the example, she is matched to it. That is she is manipulating student of $\beta^{3}$.
@@
The manipulation strategy under the constrained GS is to include an unlisted acceptable school in the list. But when the constrained GS is stable, all the seats of such a school are assigned to higher-priority students, and such manipulation does not help. This implies that constrained the serial dictatorship mechanism is non-manipulable and stable for the same set of problems.
@@
This matching is not stable at $(P,\succ,q)$ because student $i_{2}$ is unmatched, finds school $s_{3}$ acceptable while student $i_{4}$ is matched to it and $i_{2}\mathrel{\succ_{s_{3}}}i_{4}$. We claim that $GS^{2}$ is not manipulable at $(P,\succ,q)$. Only student $i_{2}$ could benefit from misrepresenting her preferences to the mechanism $GS^{2}$ because each of the other students is matched to her most preferred school. Let $P_{i_{2}}^{s_{3}}$ be a preference relation where student $i_{2}$ has ranked only school $s_{3}$ acceptable. Then,
@@
The contrasting result between the constrained Boston mechanism and the constrained Gale-Shapley can be traced back to the immediate versus deferred acceptance features and the constraint. Let us first state an important property of the Boston mechanism at the origin of its manipulability.
Lemma 1.
Let $k>1$ and $(P,\succ,q)$ a problem. Student $i$ is a manipulating student of the constrained Boston mechanism $\beta^{k}$ at the problem $(P,\succ,q)$ if and only if there is a school $s$ such that $s\mathrel{P_{i}}\beta_{i}^{k}(P,\succ,q)$, and there are less than $q_{s}$ number of students who ranked school $s$ first and have higher priority than student $i$ at school $s$.
This lemma characterizes the set of manipulating students of the Boston mechanism by those who missed schools that are not ranked first by “enough” students who have higher priority than $i$ at the school in question. Clearly, if student $i$ is a blocking student of $\beta^{k}(P,\succ,q)$, then there is a school $s$ such that $s\mathrel{P_{i}}\beta_{i}^{k}(P,\succ,q)$ and a student $j$ who is matched to school $s$ under $\beta^{k}(P,\succ,q)$ and has lower priority than $i$ at $s$ (or school $s$ has an unassigned seat under $\beta^{k}(P;\succ,q)$). Then, there are less than $q_{s}$ number of students who ranked school $s$ first and have higher priority than $i$ at $s$. Otherwise, student $j$ would not have been matched to school $s$ under $\beta^{k}(P,\succ,q)$ or no seat would have been left unassigned. This is the reason why the set of manipulating students includes blocking students.
By the deferred acceptance feature, students who are matched are neither blocking students nor manipulating students of the constrained Gale-Shapley mechanism. Blocking students and manipulating students are all students for whom the constraint is binding. That is, they are unmatched and have more acceptable schools than the ranking constraint. Because of the deferred acceptance feature, the priorities matter when a student contemplates a manipulation. No priority is violated for schools ranked within the constraint. If no student’s priority is violated in any school that she has missed, then she cannot obtain a seat at any of these schools by ranking it within the constraint. To understand this conclusion, note that for any unmatched student who replaces one acceptable school with a school where her priority is not violated, the original matching remains stable under the new problem. Since she is unmatched in a stable matching, she will be unmatched in any other stable matching of the new problem (Roth,, 1986). Then, all manipulating students are also blocking students.
@@ 5. Conclusions
In response to various concerns, many school districts around the world have recently reformed their admissions systems. The reforms are essentially two major changes. First, they replaced the immediate acceptance procedure ( Boston mechanism) with Gale-Shapley’s student-proposing deferred acceptance procedure while maintaining ranking constraints. Second, some school districts kept using the Gale-Shapley mechanism but extended the number of schools that each student is allowed to report. Anecdotal evidence points to the vulnerability to manipulation and fairness as reasons for these reforms. We showed that the immediate acceptance procedure has weakly more manipulating students than blocking students and the reverse for the Gale-Shapley’s. We demonstrated that extending ranking constraints in the Gale-Shapley mechanism led to fewer blocking students.
The fact that constrained Gale Shapley has relatively more blocking students (than it has manipulating students) suggests that in theory fairness of this mechanism might be of stronger concern than manipulability. Simultaneously, relaxing the constraint in this mechanism is guaranteed to decrease the number of blocking students. Whether this theoretical coincidence had any practical relevance for the reforms under consideration is an open question and requires further empirical research.
@@
To simplify the exposition we divide the appendix into three subsections. In each subsection, we order the results in logical order. All mechanisms that we consider are individually rational. We only consider blocking pairs to check for (the violation of) stability. We first present a useful lemma.
@@
Appendix A: Proof of Lemma 1, and Proposition 1, 3
Proof of Lemma 1.
Suppose that student $i$ is a manipulating student of $\beta^{k}$ at $(P,\succ,q)$. Then there is $P^{\prime}_{i}$ such that
(1) $\beta_{i}^{k}(P^{\prime}_{i},P_{-i},\succ,q)\mathrel{P_{i}}\beta_{i}^{k}(P,\succ,q).$
Since $\beta^{k}$ is individually rational, then $\beta_{i}^{k}(P,\succ,q)\mathrel{R_{i}}\emptyset$ and together with Equation 1, we have $\beta_{i}^{k}(P^{\prime}_{i},P_{-i},\succ,q)=s$ for some school $s\in S$. Then student $i$ did not rank school $s$ first under $P_{i}$. Suppose that there are at least $q_{s}$ students who ranked school $s$ first and have higher priority than student $i$ under $(P,\succ,q)$. Then all seats of school $s$ would have been allocated in the first step of the immediate acceptance algorithm to some students who ranked school $s$ first and have higher priority than student $i$. Then, $\beta_{i}^{k}(P^{\prime}_{i},P_{-i},\succ,q)\neq s$, which is a contradiction. Therefore, there are less than $q_{s}$ number of students who ranked school $s$ first and have higher priority than student $i$ at school $s$.
Let $(P,\succ,q)$ be a problem, $i$ a student, and $s$ a school, and suppose that $s\mathrel{P_{i}}\beta^{k}(P,\succ,q)$ and that there are less than $q_{s}$ students who ranked school $s$ first and have higher priority than student $i$ at school $s$. Thus $\beta^{k}(P_{i}^{s},P_{-i},\succ,q)=s$ and that student $i$ is manipulating student of $\beta^{k}$ at $(P,\succ,q)$. ∎
@@
Part (ii): Suppose that $\mu=FPF^{k}(P,\succ,q)$ is stable at $(P,\succ,q)$. By Claim 2 every matched student under $GS^{k}(P,\succ,q)$ is not a manipulating student of $GS^{k}$ at $(P,\succ,q)$. It is enough to show that no unmatched student under $GS^{k}(P,\succ,q)$ has a profitable misrepresentation. Because $GS^{k}$ is individually rational, by Claim 1, we further need to restrict ourselves to manipulation by top-ranking schools. Since $\mu$ is stable at $(P,\succ,q)$, we claim that it is also stable at $(P^{k},\succ,q)$. Since $GS^{k}$ is individually rational, we need to check that there is no blocking pair. Suppose, to the contrary, that a pair $(i,s)$ is a blocking pair for $\mu$ under $(P^{k},\succ,q)$. Then, $s\mathrel{P_{i}^{k}}\mu(i)$ and either (i) school $s$ has an empty seat under $\mu$ or (ii) there is a student $j$ such that $\mu(j)=s$ and $i\mathrel{\succ_{s}}j$. Note that $s\mathrel{P^{k}_{i}}\mu(i)$ implies that $s\mathrel{P_{i}}\mu(i)$. Therefore, $(i,s)$ is also a blocking pair for $\mu$ under $(P,\succ,q)$, thus contradicting our assumption that $\mu$ is stable at $(P,\succ,q)$. Therefore $\mu$ is stable at $(P^{k},\succ,q)$. Since $GS(P^{k},\succ,q)$ is the student-optimal stable matching under $(P^{k},\succ,q)$,
@@
Part (i): The Boston mechanism is a special case of the First-Preference-First mechanism when every school is a first-preference-first school. Suppose that $\beta^{k}(P,\succ,q)$ is stable at $(P,\succ,q)$. By equation 2, each student finds the outcome $GS^{k}(P,\succ,q)$ at least as good as $\beta^{k}(P,\succ,q)$ under $P^{k}$. We also know that the Boston mechanism is Pareto efficient, that is, for each problem, there is no other matching that each student finds at least as good as its outcome (Abdulkadiroğlu and Sönmez,, 2003). Therefore, the matching $\beta^{k}(P,\succ,q)=\beta(P^{k},\succ,q)$ is Pareto efficient under $P^{k}$. Thus, $GS^{k}(P,\succ,q)=\beta^{k}(P,\succ,q)$ is stable at $(P,\succ,q)$.
@@
Appendix B: Proof of Theorems 1
@@
We prove it by the contradiction. Suppose, to the contrary, that student $i$ is a blocking student for $\mu$ under $(P,\succ,q)$ such that $\mu(i)=s$ for some school $s$. Then, there is a school $s^{\prime}$ such that $s^{\prime}\mathrel{P_{i}}\mu(i)$ and either (i) $\lvert\mu^{-1}(s^{\prime})\lvert<q_{s^{\prime}}$ or (ii) there is a student $j$ such that $\mu(j)=s^{\prime}$ and $i\mathrel{\succ_{s^{\prime}}}j$. Since $\mu(i)=s$, school $s$ is one of the top $x$ acceptable schools under $P_{i}$ where $x=\ell$ if $i\in N$ and $x=k$ if $x\notin N$. Thus $s^{\prime}\mathrel{P^{x}_{i}}\mu(i)=s$ and $(i,s^{\prime})$ is a blocking pair of $\mu$ under $(P^{\ell}_{N},P^{k}_{-N},\succ,q)$, contradicting the stability of $\mu$ under $(P^{\ell}_{N},P^{k}_{-N},\succ,q)$.
@@
We call on to the sequential version of McVitie and Wilson, (1970) of the deferred acceptance algorithm. This is a version where students apply one at a time according to a predetermined order such that in each step the highest-ordered student among the ones whose applications have not yet been tentatively accepted applies.
The idea of the proof is to consider close ranking constraints $k-1$ and $k$, where $k>1$, and replace students’ preference relations in $P^{k-1}$ with the ones in $P^{k}$. Let $N\subsetneq I$ be a proper subset of $I$ and $i\notin N$. Suppose that starting from $P^{k-1}$ we have replaced all the preferences of students in $N$ by their preferences in $P^{k}$ and define $\widehat{P}=(P_{N}^{k},P^{k-1}_{-N})$. Note that $N$ may be empty. Next we consider student $i$: from $\widehat{P}$ we replace her preference relation $P^{k-1}_{i}$ by $P^{k}_{i}$. Let $\nu=GS(P_{i}^{k-1},\widehat{P}_{-i},\succ,q)$ and $\mu=GS(P^{k}_{i},\widehat{P}_{-i},\succ,q)$.